Write True (T) or False ( F ) against each of the following statements : 
0.9 : 0.36 :: 10 : 4

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

0.9: 0.36:: 10: 4
0.9 / 0.36 = 90 / 36
= 10 / 4
Hence, 0.9: 0.36 = 10: 4
Therefore True

Thanks